OpenStudy (anonymous):

The line whose equation is y= 4x+2 has a y-intercept whose coordinates are 1) (0,0) 2) (0,2) 3) (4,0) 4) (0,4) PLEASE HELP!

OpenStudy (amistre64):

-4x and divide by 2 -4x +y = 2 -4x y --- + --- = 1 2 2 -2x y --- + --- = 1 ; this is called the intercept form of a line 1 2

OpenStudy (amistre64):

when x=0; y=2 when y=0;x=-1/2

OpenStudy (anonymous):

?

OpenStudy (amistre64):

y intercept is the x=0 axis; (0,2)

OpenStudy (anonymous):

y intercept set x = 0 and solve for y with your eyeballs. y = 2

OpenStudy (amistre64):

the intercept form doesnt get alot of publicity does it :) http://www.mathwords.com/t/two_intercept_form_for_the_equation_of_a_line.htm

OpenStudy (amistre64):

it prolly more for thos: ax +by = c equations

OpenStudy (anonymous):

@amistre64 THANKS YOU!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

=)